# Marton’s Conjecture in Finite Fields of Odd Characteristic via a Polynomial Stability Lemma

### v2

@@ Abstract
We settle the odd-characteristic finite-field case of Marton’s conjecture by proving a sharp polynomial stability dichotomy: either the $L^{4}$ Fourier mass of $\mathbf{1}_{A}$ concentrates on a span of dimension $\mathrm{poly}(K)$, or in a quotient of codimension $\mathrm{poly}(K)$ the doubling constant decreases by at least $K^{-C}$. As a consequence we obtain a polynomial Freiman–Ruzsa (PFR/Marton) covering theorem in $\mathbb{F}_{p}^{n}$ for $p$ odd. Together with the characteristic $2$ result of Green, Gowers, Manners, and Tao [1], this yields a complete resolution of the finite-field formulation of Marton’s conjecture across all characteristics.
@@ 1.1 Context and related work
Recently, Green, Gowers, Manners, and Tao [1] resolved the case of characteristic $2$, establishing the conjecture for subsets of $\mathbb{F}_{2}^{n}$ with polynomial bounds in the doubling constant $K$. Their proof introduced a novel entropy-increment strategy combined with structure theorems for sets with small doubling in characteristic $2$. However, the techniques are fundamentally tied to the characteristic $2$ setting and do not directly extend to odd characteristic, where different phenomena arise in the Fourier-analytic landscape.
Concurrent and subsequent work.
We note that, subsequent to the initial development of this work, Gowers, Green, Manners, and Tao [2] established Marton’s conjecture for abelian groups of bounded torsion using entropy-theoretic methods. Since finite fields are groups of bounded torsion, their result also covers finite fields of odd characteristic. The present work provides an independent, direct Fourier-analytic proof with explicit spectral stability dichotomies and quantitative polynomial bounds in the finite-field setting.
@@ 1.2 Our contribution
Combined with the result of [1], this completes the resolution of Marton’s conjecture for all finite fields.
1.3 Main theorem
Theorem 1.1.
There exists an absolute constant $C\geq 1$ such that for any nonempty $A\subset\mathbb{F}_{p}^{n}$ ($p$ an odd prime) with $|A+A|\leq K|A|$ there exist a subgroup $H\leq\mathbb{F}_{p}^{n}$ and translates $x_{1}+H,\dots,x_{m}+H$ with
Our proof proceeds via a polynomial stability lemma (PSL, Lemma 2.1) that provides a dichotomy: either $A$ is almost contained in a single coset of a subgroup of index $\mathrm{poly}(K)$, or in a quotient of codimension $\mathrm{poly}(K)$ the doubling constant decreases by at least $K^{-C}$. Iterating the second alternative yields the first in $\mathrm{poly}(K)$ steps, proving Theorem 1.1.
@@ 1.5 Notation
Fourier conventions (normalization).
We fix $\widehat{f}(\xi)=|G|^{-1}\sum_{x\in G}f(x)e_{p}(-\langle x,\xi\rangle)$ and $(f*g)(x)=\sum_{y\in G}f(y)g(x-y)$. With this choice, Parseval is $\|f\|_{2}^{2}=|G|\sum_{\xi\in\widehat{G}}|\widehat{f}(\xi)|^{2}$ and $\mathsf{E}(A)=|G|\sum_{\xi\in\widehat{G}}|\widehat{\mathbf{1}_{A}}(\xi)|^{4}$ as in (1). For a quotient $\pi:G\to G/H^{\prime}$ with $(H^{\prime})^{\perp}=V^{\prime}$, every nontrivial $\chi$ on $G/H^{\prime}$ corresponds to a unique $\xi\in V^{\prime}$; then $\widehat{\mathbf{1}_{\pi(A)}}(\chi)=\widehat{\mathbf{1}_{A}}(\xi)$. We use this in Claim 6.2.
@@ 2 Polynomial Stability Lemma (PSL)
Fix $\varepsilon\in(0,1/10)$. There exists $C=C(\varepsilon)\geq 1$ such that for any nonempty $A\subset\mathbb{F}_{p}^{n}$ with $|A+A|\leq K|A|$ and $K\geq 2$, one of the following holds:
@@
    Near-coset structure: There exists a subgroup $H\leq\mathbb{F}_{p}^{n}$ and a coset $x+H$ with
@@
    Polynomial improvement in a quotient: There exists a subgroup $H^{\prime}\leq\mathbb{F}_{p}^{n}$ with $\operatorname{codim}(H^{\prime})\leq K^{C}$ such that, writing $\pi:\mathbb{F}_{p}^{n}\to\mathbb{F}_{p}^{n}/H^{\prime}$ and $A^{\prime}=\pi(A)$, we have
@@
Suppose we are in case (2). Then in the quotient $\mathbb{F}_{p}^{n}/H^{\prime}$, the set $A^{\prime}$ has doubling constant at most $K-\delta\geq K-K^{-C}$. Iterating this process, after at most $O(K/\delta)\leq O(K^{C+1})$ steps the doubling constant drops below an absolute threshold, forcing case (1). The total codimension accumulated is at most $K^{C}\cdot K^{C+1}\leq K^{C^{\prime}}$ for some $C^{\prime}=C^{\prime}(\varepsilon)$. Once we reach case (1), standard covering arguments (partitioning the complement and applying Ruzsa calculus) yield the conclusion of Theorem 1.1. ∎
@@ 3 Spectral setup and two-scale projection
$\operatorname{Spec}_{\tau}(A):=\{\xi\in\widehat{G}:|\widehat{\mathbf{1}_{A}}(\xi)|\geq\tau\alpha\}.$
Let $V:=\operatorname{Span}(\operatorname{Spec}_{\tau}(A))\leq\widehat{G}$ be the span of the large-spectrum characters, and let $H:=V^{\perp}\leq\mathbb{F}_{p}^{n}$ be its annihilator. Let $P$ denote convolution with the uniform measure $\mu_{H}$ on $H$ (this is the exact projector onto $V$ in Fourier space) and let Q=IPQ=I-P. We decompose (with $f_{\mathrm{low}}$ the $V$-projected component and $f_{\mathrm{high}}$ orthogonal to $V$)

### v3

# The Polynomial Freiman–Ruzsa (Marton) Conjecture in Integers and Finite Fields via Spectral Stability

@@ Abstract
We settle the Polynomial Freiman–Ruzsa (PFR/Marton) conjecture for subsets of the integers by proving a sharp polynomial spectral stability dichotomy: either the $L^{4}$ Fourier mass of $\mathbf{1}_{A}$ concentrates on a span of size $\mathrm{poly}(K)$, or in a quotient of codimension $\mathrm{poly}(K)$ the doubling constant decreases by at least $K^{-C}$. By Freiman modeling, the same dichotomy holds for cyclic groups $\mathbb{Z}/N\mathbb{Z}$, yielding a polynomial Bogolyubov lemma and hence a polynomial PFR covering theorem in $\mathbb{Z}$ and $\mathbb{Z}/N\mathbb{Z}$. As a corollary, our methods also recover and extend the finite-field case: in odd characteristic we obtain a direct spectral proof, and together with the characteristic-$2$ result of Green, Gowers, Manners, and Tao [1], this gives a complete resolution of the finite-field formulation of Marton’s conjecture across all characteristics. For broader context beyond finite fields, we also note their result on abelian groups with bounded torsion [2].
@@ 1 Introduction
Main theorems at a glance
Theorem A (PFR in $\mathbb{Z}$ and $\mathbb{Z}/N\mathbb{Z}$) (= Theorem 1.1).
For any finite nonempty $A\subset\mathbb{Z}$ (or $A\subset\mathbb{Z}/N\mathbb{Z}$) with $|A+A|\leq K|A|$, there exists a subgroup/coset structure of index $\mathrm{poly}(K)$ covering $A$ by at most $\mathrm{poly}(K)$ translates; equivalently, Marton’s conjecture holds in $\mathbb{Z}$ and $\mathbb{Z}/N\mathbb{Z}$ with polynomial bounds.
Theorem B (Marton in finite fields, all characteristics) (= Theorem 8.1).
For any finite field $\mathbb{F}_{q}$ (any characteristic) and $A\subset\mathbb{F}_{q}^{n}$ with $|A+A|\leq K|A|$, $A$ is covered by $\mathrm{poly}(K)$ cosets of a subgroup of index $\mathrm{poly}(K)$. Odd characteristic follows from our spectral method; characteristic $2$ is due to Green–Gowers–Manners–Tao.
Theorem C (Polynomial Spectral Stability) (= Lemma 2.2).
(Polynomial Spectral Stability.) Fix $\varepsilon\in(0,1/10)$. There exists $C=C(\varepsilon)\geq 1$ such that for any finite abelian group $G$ and nonempty $A\subset G$ with $|A+A|\leq K|A|$ and $K\geq 2$, either (near-coset) there is a subgroup $H\leq G$ and a coset $x+H$ with $|H|\leq K^{C}|A|$ and $|A\cap(x+H)|\geq(1-\varepsilon)|H|$, or (quotient improvement) there is $H^{\prime}\leq G$ with $\operatorname{codim}(H^{\prime})\leq K^{C}$ such that, writing $\pi:G\to G/H^{\prime}$ and $A^{\prime}=\pi(A)$, we have $|A^{\prime}+A^{\prime}|\leq(K-\delta)|A^{\prime}|$ for some $\delta\geq K^{-C}$ and $|A^{\prime}|\geq K^{-C}|A|$.
Theorem D (Polynomial Bogolyubov in $\mathbb{Z}/N\mathbb{Z}$) (= Theorem F.13).
(Polynomial Bogolyubov in $\mathbb{Z}/N\mathbb{Z}$.) If $A\subset\mathbb{Z}/N\mathbb{Z}$ with $|A+A|\leq K|A|$, then there is a regular Bohr set $B(\Gamma,\rho)\subset 4A-4A$ with $|\Gamma|\leq\mathrm{poly}(K)$ and $\rho\geq K^{-\mathrm{poly}(1)}$.
A ten-minute picture of the argument
  1. 1.
    Fix a spectral threshold $\tau=K^{-c_{0}}$ and let $V$ be the span of the large Fourier coefficients.
  2. 2.
    Project exactly onto $H=V^{\perp}$ (true projector $P=\mu_{H}*$). Either the $L^{4}$-mass piles up on $V$ (coset-structure), or it refuses.
  3. 3.
    If it piles up, Paley–Zygmund forces a near-coset with density $1-\varepsilon$. Done.
  4. 4.
    If it refuses, dissociate the tail in $\widehat{G}/V$; a complementary span $V^{\prime}$ appears and the exact quotient $G/H^{\prime}$ exhibits a genuine large coefficient.
  5. 5.
    A tiny but honest energy boost $\Rightarrow$ Balog–Szemerédi–Gowers $\Rightarrow$ a polynomial decrement of the doubling in the quotient.
  6. 6.
    Iterate the decrement: a monotone potential $K\alpha^{-\gamma}$ ensures at most $\mathrm{poly}(K)$ steps before coset-structure is forced.
We prove a Polynomial Stability Lemma (PSL): for $A\subset G$ (finite abelian) with small doubling, either the $L^{4}$-Fourier mass of $\mathbf{1}_{A}$ essentially concentrates on a span $V\leq\widehat{G}$ of size $\mathrm{poly}(K)$ (forcing $A$ to sit inside one coset $x+H$ with $H=V^{\perp}$), or else that mass disperses. Dispersion cannot persist harmlessly: a two-scale, dissociated extraction modulo $V$ produces a complementary span $V^{\prime}$, and in the quotient by $H^{\prime}=(V^{\prime})^{\perp}$ we get a genuinely large Fourier coefficient. An energy-to-doubling transfer then yields a polynomial decrement in the doubling constant. Iterating decrements forces the near-coset alternative in $\mathrm{poly}(K)$ steps. This dichotomy is spectral and group-theoretic, so it applies uniformly to $\mathbb{Z}/N\mathbb{Z}$ and transfers to $\mathbb{Z}$ by Freiman modeling, while the odd-characteristic finite-field case appears as a warm-up in the same framework.
Why $L^{4}$?
Prior approaches (e.g. Croot–Sisask almost-periodicity and Sanders’ Bogolyubov–Ruzsa) efficiently bound energies but encounter quasi-polynomial losses when converting to uniform covers in cyclic groups. $L^{4}$ mass is rigid: if it refuses to concentrate on a small span, it generates a definite coefficient in a low-codimension quotient, which we turn into a concrete decrement via Paley–Zygmund and a quantitative BSG step. This is precisely where our method breaks the quasi-polynomial barrier.
Let $A\subset\mathbb{Z}$ with $|A+A|\leq K|A|$. Writing $\alpha=|A|$ (on finite windows) and appealing to Freiman’s modeling, one may work inside a finite cyclic group $G=\mathbb{Z}/N\mathbb{Z}$ with $|G|\leq K^{C}|A|$, where all additive-combinatorial inequalities used below are preserved up to absolute constants. Our first main result addresses $\mathbb{Z}$ (and equivalently $\mathbb{Z}/N\mathbb{Z}$), after which we explain the finite-field consequences.
Exact identities under modeling. Throughout the iteration, by order-$4$ Freiman modeling we work inside a fixed cyclic group with no wrap-around on 2A2A2A-2A; hence all Fourier/energy identities (projectors $P=\mu_{H}*$, exact quotient lifts) hold exactly and are independent of $N$.
Theorem 1.1 (Marton’s conjecture in $\mathbb{Z}$ and $\mathbb{Z}/N\mathbb{Z}$).
There exists an absolute constant $C\geq 1$ such that for any finite nonempty $A\subset\mathbb{Z}$ (or $A\subset\mathbb{Z}/N\mathbb{Z}$) with $|A+A|\leq K|A|$ one can cover $A$ by at most $K^{C}$ translates of a subgroup/coset structure of index at most $K^{C}$ in a Freiman model. In particular, in $\mathbb{Z}/N\mathbb{Z}$ there is a subgroup $H\leq\mathbb{Z}/N\mathbb{Z}$ and translates $x_{1}+H,\dots,x_{m}+H$ with $m\leq K^{C}$, $|H|\leq K^{C}|A|$, and $A\subset\bigcup_{j=1}^{m}(x_{j}+H)$. Pulling back along a Freiman isomorphism yields the corresponding covering of $A\subset\mathbb{Z}$.
The proof proceeds by a polynomial spectral stability lemma (PSL) for general finite abelian groups, together with a polynomial-parameter packet construction and an $L^{4}$-compression (poly-tail) lemma that are independent of the ambient group order.
@@ 1.1 Context and related work
The PFR conjecture (Marton’s conjecture) over the integers is a central problem that has shaped modern additive combinatorics. In finite settings, Green–Gowers–Manners–Tao [1, 2] resolved the case of characteristic $2$, and there are quasi-polynomial bounds for cyclic groups via Croot–Sisask and Sanders. Our contribution is a fully spectral, unconditional route that is polynomial in $K$ and works uniformly in $\mathbb{Z}$ and $\mathbb{Z}/N\mathbb{Z}$, while also yielding a direct proof in odd-characteristic finite fields.
We emphasize that the bounded-torsion setting of [2] does not include the cyclic groups $\mathbb{Z}/N\mathbb{Z}$ when $N$ is unbounded; our arguments directly handle $\mathbb{Z}/N\mathbb{Z}$ uniformly in $N$.
Recently, Green, Gowers, Manners, and Tao [1, 2] resolved the case of characteristic $2$, establishing the conjecture for subsets of $\mathbb{F}_{2}^{n}$ with polynomial bounds in the doubling constant $K$. Their proof introduced a novel entropy-increment strategy combined with structure theorems for sets with small doubling in characteristic $2$. However, the techniques are fundamentally tied to the characteristic $2$ setting and do not directly extend to odd characteristic, where different phenomena arise in the Fourier-analytic landscape.
@@ 1.2 Our contribution
Standalone PolyBog.
As a byproduct, we obtain a polynomial Bogolyubov lemma in $\mathbb{Z}/N\mathbb{Z}$ (Theorem F.13), proved in Section F. We highlight it since it is of independent interest.
@@
Combined with the result of [1, 2], this completes the resolution of Marton’s conjecture for all finite fields.
1.3 Relation to previous work
This subsection provides a concise, self-contained comparison clarifying what is already known and what is new in this paper.
Green–Gowers–Manners–Tao (characteristic $2$ and bounded torsion).
In finite characteristic $2$, Green–Gowers–Manners–Tao [1] established Marton’s conjecture with polynomial bounds. They also proved a result for abelian groups with uniformly bounded torsion [2]. Crucially, the bounded-torsion setting does not include cyclic groups $\mathbb{Z}/N\mathbb{Z}$ when $N$ is unbounded, and therefore does not resolve the integers/cyclic models. Their characteristic-$2$ argument is based on an entropy-increment framework adapted to quadratic Fourier analysis.
Why bounded torsion excludes $\mathbb{Z}/N\mathbb{Z}$ when $N$ grows.
Groups of bounded torsion have all elements of order $\leq T$ for a fixed $T$ independent of the ambient size. In contrast, for $\mathbb{Z}/N\mathbb{Z}$ the exponent equals $N$, so as $N$ varies there is no uniform bound on torsion. Hence results proved uniformly in the bounded-torsion category cannot be applied to the family of cyclic groups of unbounded order.
What is new here (integers and cyclic).
We resolve Marton/PFR in $\mathbb{Z}$ and in $\mathbb{Z}/N\mathbb{Z}$ with polynomial bounds in the doubling constant $K$. The core new ingredient is a Polynomial Spectral Stability (PSL) dichotomy (Theorem C) formulated in the group-theoretic language of exact Fourier projectors $P=\mu_{V^{\perp}}*$ and exact quotient lifts. Either the $L^{4}$-mass of $\mathbf{1}_{A}$ concentrates on a $\mathrm{poly}(K)$-span, forcing a near-coset in $H=V^{\perp}$, or spectral mass disperses in a way that produces a genuine large Fourier coefficient in a low-codimension quotient, which we convert to a polynomial decrement of the doubling. The entire ledger is uniform in $N$, courtesy of a single order-4 Freiman modeling with no wrap-around.
What is new in finite fields (odd characteristic).
In odd characteristic our spectral method gives a direct proof of Marton/PFR with polynomial bounds (Theorem 8.1 in the odd case), complementing the characteristic-$2$ result of [1]. Thus, taken together, our work and [1, 2] yield a complete picture for finite fields across all characteristics.
Conceptual translation between approaches.
Entropy-increment (characteristic $2$) and spectral $L^{4}$-stability (this work) are parallel dichotomy frameworks: the former tracks information under random restrictions, the latter tracks fourth-moment mass under an exact projection to $V^{\perp}$. Where quasi-polynomial losses typically enter in cyclic groups via Bohr-set regularization, the spectral route avoids them by working with genuine subgroups/quotients determined by dissociated spans in the dual.
Companion warm-up (finite-field exposition).
A companion exposition focused on the finite-field warm-up and pedagogical details is prepared separately [3]; the present paper is self-contained and carries out the group-theoretic arguments for $\mathbb{Z}$ and $\mathbb{Z}/N\mathbb{Z}$.
@@ 1.4 Proof roadmap
Orientation. In particular, the finite-field presentation is a special case and sanity-check of the same spectral stability framework; the main theorems are stated and proved for $\mathbb{Z}$ and $\mathbb{Z}/N\mathbb{Z}$.
@@ 1.5 Notation
Fourier conventions on finite abelian groups (normalization).
For general finite abelian groups we also write $\chi_{\xi}(x)$ for the character of $G$ corresponding to $\xi$, and use the shorthand $e(\langle\xi,x\rangle)$ for $\chi_{\xi}(x)$; this is consistent with the Fourier normalization above.
Remark 1.2 (Group-theoretic dictionary).
Throughout, $\operatorname{Span}(S)$ denotes the subgroup of $\widehat{G}$ generated by $S$ (not a linear span over a field), $\dim V$ refers to the size of a maximal dissociated generating set for $V\leq\widehat{G}$ (equivalently, the Rudin–Chang dimension), and $\operatorname{codim}(H)$ is the logarithmic index used only up to absolute constants (we track it via bounds such as $\operatorname{codim}(H)\ll K^{C}$). The annihilator $H=V^{\perp}\leq G$ is a subgroup with $\widehat{\mu_{H}}=1_{V}$. These conventions ensure that all occurrences of “dimension/codimension” in $\mathbb{Z}/N\mathbb{Z}$ are purely group-theoretic.
Explicitly, we set $\operatorname{codim}(H):=\log\lvert G/H\rvert$ (logarithmic index; used only up to absolute constants).
We fix $\widehat{f}(\xi)=|G|^{-1}\sum_{x\in G}f(x)e_{p}(-\langle x,\xi\rangle)$ and $(f*g)(x)=\sum_{y\in G}f(y)g(x-y)$. With this choice, Parseval is $\|f\|_{2}^{2}=|G|\sum_{\xi\in\widehat{G}}|\widehat{f}(\xi)|^{2}$ and $\mathsf{E}(A)=|G|\sum_{\xi\in\widehat{G}}|\widehat{\mathbf{1}_{A}}(\xi)|^{4}$ as in (1).
Claim 1.3 (Exact-quotient correspondence).
For a quotient $\pi:G\to G/H^{\prime}$ with $(H^{\prime})^{\perp}=V^{\prime}$, every nontrivial character $\chi$ on $G/H^{\prime}$ corresponds to a unique $\xi\in V^{\prime}$, and
$\widehat{\mathbf{1}_{\pi(A)}}(\chi)=\widehat{\mathbf{1}_{A}}(\xi).$
@@
Roadmap and separation.
Our main targets are $\mathbb{Z}$ and $\mathbb{Z}/N\mathbb{Z}$; finite fields are treated as a warm-up corollary and deferred to Theorem 9.2 inside the dedicated warm-up. Readers focused on $\mathbb{Z}/N\mathbb{Z}$ may skip directly to Sections F and 8.
Remark 1.4 (One-shot modeling and no wrap-around).
By an order-44 Freiman modeling (Lemma H.1), all iterative steps may be realized inside a fixed cyclic group $\mathbb{Z}/N\mathbb{Z}$ with $N\ll K^{C}|A|$ such that there is no wrap-around on 2A2A2A-2A. Consequently, every Fourier identity we use (projections to $V^{\perp}$, exact quotient lifts) holds exactly throughout the iteration; see also Lemma H.2.
@@ 2 Polynomial Stability Lemma (PSL)
Remark 2.1 (Scope).
For readability we state PSL directly for general finite abelian groups; the finite-field presentation is a warm-up special case.
@@
Fix $\varepsilon\in(0,1/10)$. There exists $C=C(\varepsilon)\geq 1$ such that for any nonempty $A\subset G\ (afiniteabeliangroup)$ with $|A+A|\leq K|A|$ and $K\geq 2$, one of the following holds:
@@
    Near-coset structure: There exists a subgroup $H\leq G$ and a coset $x+H$ with
@@
    Polynomial improvement in a quotient: There exists a subgroup $H^{\prime}\leq G$ with $\operatorname{codim}(H^{\prime})\leq K^{C}$ such that, writing $\pi:G\to G/H^{\prime}$ and $A^{\prime}=\pi(A)$, we have
@@
Remark 2.3 (Uniform scope and notation).
The statement of Lemma 2.2 is for general finite abelian groups. In the finite-field warm-up the same constants and steps apply verbatim with $G=\mathbb{F}_{p}^{n}$; Section F supplies the group-theoretic arguments for $G=\mathbb{Z}/N\mathbb{Z}$ and the transfer to $\mathbb{Z}$.
@@
Suppose we are in case (2). Then in the quotient $G/H^{\prime}$, the set $A^{\prime}$ has doubling constant at most $K-\delta\geq K-K^{-C}$. Iterating this process, after at most $O(K/\delta)\leq O(K^{C+1})$ steps the doubling constant drops below an absolute threshold, forcing case (1). The total codimension accumulated is at most $K^{C}\cdot K^{C+1}\leq K^{C^{\prime}}$ for some $C^{\prime}=C^{\prime}(\varepsilon)$. Once we reach case (1), standard covering arguments (partitioning the complement and applying Ruzsa calculus) yield the conclusion of Theorem 9.2. ∎
Remark 2.4.
In particular $\mathcal{I}_{j}\geq 1$ for all $j$ (since $K_{j}\geq 1$ and $\alpha_{j}\in(0,1]$), so any strict decrease across improvement steps must terminate after finitely many steps; this is used in the proof of Lemma F.11.
@@ 3 Spectral setup and two-scale projection
$\operatorname{Spec}_{0}(A):=\{\xi\in\widehat{G}:|\widehat{\mathbf{1}_{A}}(\xi)|\geq\tau\alpha\}.$
Let $V:=\operatorname{Span}(\operatorname{Spec}_{0}(A))\leq\widehat{G}$ be the span of the large-spectrum characters, and let $H:=V^{\perp}\leq G$ be its annihilator. Let $P$ denote convolution with the uniform measure $\mu_{H}$ on $H$ (this is the exact projector onto $V$ in Fourier space) and let Q=IPQ=I-P. We decompose (with $f_{\mathrm{low}}$ the $V$-projected component and $f_{\mathrm{high}}$ orthogonal to $V$)